# A Survey on LLM-as-a-Judge

### v1

@@ 5.1.2. Evaluation Metrics
For position bias, we use Position Consistency as metric, which quantifies how frequently a judge model selects the same response after changing their positions. Formally, given $N$ samples $\{(q_{i},r1_{i},r2_{i})\}^{N}_{i=1}$, for each sample $(q_{i},r1_{i},r2_{i})$, we query the LLM evaluator with two prompts $P(q_{i},r1_{i},r2_{i})$ and $P(q_{i},r2_{i},r1_{i})$, and obtain corresponding two evaluation results $S_{i}^{r12}$ and $S_{i}^{r21}$. Each $S_{i}$ is $r1_{i}$, $r2_{i}$ or ”TIE”. Then we calculate the position consistency as follows:
@@ 5.1.3. Target LLMs and Strategies
For LLMs, we selected three LLMs commonly used in automatic evaluation as evaluators, including closed-source LLMs GPT-4, GPT-3.5, and open-source LLMs LLaMA3. For improvement strategies, we selected Providing Evaluations with Explanations, Self Validation, Summarize by Multiple Rounds, and Vote by Multiple LLMs, since these strategies are all very straightforward and relatively uniform in many works. We adopt GPT-3.5 as the base evaluator when conducting meta-evaluation for these improvement strategies.
@@ 5.1.4. Model Configuration
For the closed-source LLMs, we use the official API provided by OpenAI to interact with LLMs. The model versions we selected are GPT-4-turbo and GPT-3.5-turbo, specifically referencing gpt-4-turbo-2024-04-09 and gpt-3.5-turbo-0125 respectively [^2] https://platform.openai.com/docs/models. For the open-source LLM, we adopt Meta-Llama-3-8B-Instruct [^3] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct provided by meta-llama on Hugging Face, deployed on an Ubuntu machine equipped with a 40GB NVIDIA A100 GPU. To stabilize the evaluation results of LLMs, we set the hyper parameter temperature to 0 to reduce the impact of randomness in LLMs’ output. For Summarize by Multiple Rounds, we conduct 5 rounds for each sample.
@@ 5.2. Experiment Results and Analysis
Model Biases Alignment Position Length Concre- Empty Content Nested Familiar with Human teness Reference Continuation Instruction Knowledge (n=5106) (n=2633) (n=34) (n=28) (n=26) (n=24) (n=24) (n=24) GPT-4-turbo 61.57 80.49 91.18 89.29 65.38 95.83 70.83 100.0 LLaMA3-8B-Instruct 50.72 38.85 20.59 57.14 65.38 75.00 45.83 54.17 GPT-3.5-turbo 54.72 68.78 20.59 64.29 23.08 91.67 58.33 54.17 GPT-3.5-turbo - w/ explanation 52.47 48.97 35.29 60.71 38.46 91.67 41.67 50.00 - w/ self-validation 54.86 69.31 23.53 60.71 23.08 91.67 41.67 50.00 - w/ multi rounds 54.68 70 11 26.47 67.86 23.08 95.83 54.17 50.00 Multi LLMs 57.66 32.28 26.47 64.28 46.15 87.50 66.67 62.50
@@
Table 2. The meta-evaluation results for LLMs and improvement strategies based on GPT-3.5-turbo. All the values are percentages.
The experiment results are shown in Table 2. Comparing the evaluation performance of different LLMs, we found significant differences between them. GPT-4 outperforms GPT-3.5 and LLaMA3-8B-Instruct with a large margin across all meta-evaluation dimensions and shows fewer biases. Therefore, when economically feasible, using GPT-4 for automatic evaluation may obtain objective results with minimal biases. LLaMA3-8B-Instruct and GPT-3.5 have similar metrics across most meta-evaluation dimensions. However, LLaMA3-8B-Instruct performs poorly in position bias, while GPT-3.5 struggles in empty reference bias. Since LLaMA3-8B-Instruct is open-source, it may also achieve good evaluation performance as a relatively lightweight automatic evaluator after specific debiasing fine-tuning.
@@
Comparing the performance of different improvement strategies, it reveals that no all modifications effectively improve LLM-as-a-judge’s evaluation outcomes. Providing with Explanation (- w/ explanation) provides interpretability by offering reasons alongside evaluation scores or selections, which aids in logical backtracking during human review. However, in terms of evaluation performance and bias mitigation, it generally has a negative impact. This performance decline is speculated to be caused by deeper biases introduced by self-explanation. Self Validation (- w/ self-validation) shows minimal effectiveness, likely due to the LLMs’ overconfidence, which may limit its re-evaluation efforts during self-validation. We will further discuss this limitation in Section 7.1.
Summarize by Multiple Rounds (- w/ multi rounds) is a strategy with clear benefits, showing improvements across multiple dimensions. It suggests that repeated evaluations help reduce the impact of randomness in LLMs, thereby addressing bias issues. Vote by Multiple LLMs (Multi LLMs) shows a noticeable decline in performance of position bias, but with either slight improvements or minimal differences in other dimensions, when compared to GPT-3.5-turbo. This trend is influenced by the comprehensive performance of GPT-4-turbo, GPT-3.5-turbo, and LLaMA3-8B-Instruct. GPT-4-turbo generally aids in correction, but it is significantly impacted in position bias, where LLaMA3-8B-Instruct underperforms. Similarly, GPT-3.5-turbo’s weak performance in handling empty reference bias notably reduces the corresponding metric. This indicates that when multiple LLMs are used for joint evaluation, the differences in their capabilities need to be carefully considered.
@@
In summary, due to the inherent capabilities and potential risks of LLMs, common improvement strategies for LLM-as-a-judge are not fully effective in improving the evaluation performance or mitigating biases. The limitations and challenges will be further discussed in Section 7.
@@ 6.2.1. Finance
LLMs have demonstrated significant potential in the finance domain, particularly in tasks such as forecasting, anomaly detection, and personalized text generation (Zhao et al., 2024), thereby driving an increasing demand for LLM evaluators.
In the context of LLM-as-a-judge applications within finance, expert knowledge is crucial for domain-specific evaluations. Current research can be divided into two areas: one focuses on designing LLM-based evaluators that leverage expert knowledge for specific tasks. For instance, Brief et al. (2024) conducted a case study on multi-task fine-tuning in finance to enhance LLM performance (Brief et al., 2024), while Yu et al. (2024) introduced FinCon, a multi-agent system that uses conceptual verbal reinforcement to improve financial decision-making (Yu et al., 2024). The second area of research aims to provide benchmarks to evaluate and enhance LLMs’ understanding of domain-specific knowledge. These benchmarks include user-feedback-based UCFE (Yang et al., 2024), IndoCareer—a dataset of professional exam questions (Koto, 2024), and AI-generated domain-specific evaluation sets (Raju et al., 2024).
Additionally, the concept of LLM-as-a-judge shows promising applications in credit scoring (Yoon, 2023; Babaei and Giudici, 2024) and Environmental, Social, and Governance (ESG) scoring (Zhao et al., 2024). This work remains in its early stages, necessitating further exploration to refine evaluation methods and expand applications in the finance domain.
@@ 6.2.2. Law
The second category of research involves creating benchmarks for evaluating LLM applicability in legal scenarios. Examples include multi-domain evaluation sets, such as the IndoCareer dataset for professional exams in Indonesia (Koto, 2024) and LegalBench, a collaboratively built benchmark for assessing legal reasoning capabilities in LLMs across multiple domains and languages (Guha et al., 2024). These benchmarks are often language-specific like LexEval for Chinese legal texts (Li et al., 2024a) and Eval-RAG for Korean  (Ryu et al., 2023). Other benchmarks target specific attributes, such as ethics (Zhang et al., 2024) and harmfulness (Andriushchenko et al., 2024).
@@ 6.2.4. Others
LLMs have been employed as evaluators across various fields to enhance efficiency and consistency. In software engineering, a method was proposed for using LLMs to evaluate bug report summarizations, demonstrating high accuracy in assessing correctness and completeness, even surpassing human evaluators who experienced fatigue. This approach offers a scalable solution for evaluation (Kumar et al., 2024a). In education, automated essay scoring and revising were explored using open-source LLMs, achieving performance comparable to traditional deep-learning models. Techniques such as few-shot learning and prompt tuning improved scoring accuracy, while revisions effectively enhanced essay quality without compromising original meaning (Song et al., 2024). In content moderation, an LLM-based approach was developed to identify rule violations on platforms like Reddit, achieving high true-negative rates but encountering challenges with complex rule interpretation, emphasizing the necessity of human oversight for nuanced cases (Kolla et al., 2024). These applications of LLMs as evaluators highlight their growing potential in diverse sectors, emphasizing the need for integrating domain-specific knowledge and refining methodologies.
@@ 7.1. Reliability
Evaluating the reliability of LLMs when used as judges reveals several pressing challenges. Both human and LLM judges exhibit biases, which raises concerns regarding the consistency and fairness of their evaluations. Specifically, human judges are also found to have inherent bias (Zheng et al., 2023b; Wu and Aji, 2023) and may not even provide reliable answers (Clark et al., 2021; Hämäläinen et al., 2023). As an alternative to humans, LLM evaluations are also found to have certain biases, and the annotation results require more evaluation (Pangakis et al., 2023), as we discussed in § 4. The bias of LLM-as-a-judge is more due to the fact that LLM is a probabilistic model, as we have defined in § 4. Moreover, Reinforcement Learning with Human Feedback (RLHF) improves LLM performance by aligning them with human preferences. However, ensuring models trained with RLHF produce robust and consistent outputs remains an ongoing challenge.
@@
Instruction-tuned LLMs have been demonstrated to possess the issue of overconfidence, which means they tend to offer overly favorable scores when evaluating their own responses (Tian et al., 2023a). The overconfidence is also highly likely to exist in the scenario of LLM-as-a-judge, which is also engaged in evaluating the responses generated by LLMs. Consequently, when LLM-as-a-judge is utilized with the latest LLMs, which are typically instruction-tuned, the existence and impact of overconfidence need to be meticulously examined.
@@ 8.1. More Reliable LLM-as-a-Judge
Finally, the uncertain and evolving nature of robustness risks underscores the necessity of proactive mitigation strategies. These strategies should include the development of adversarial training techniques tailored to judgment tasks, the integration of robust uncertainty quantification methods, and the implementation of human-in-the-loop systems to oversee critical decisions. By addressing these challenges, we can build more resilient and dependable systems capable of maintaining high levels of reliability even under adversarial conditions.
@@ 8.2. LLM-as-a-Judge for Data Annotation
Despite its wide applications, data annotation poses significant challenges for current machine learning models due to the complexity, subjectivity, and diversity of data. This process requires domain expertise and is resource-intensive, particularly when manually labeling large datasets. Advanced LLMs such as GPT-4 (OpenAI, 2023a), Gemini (Google, 2023), and LLaMA-2 (Touvron et al., 2023c) offer a promising opportunity to revolutionize data annotation. LLMs serve as more than just tools but play a crucial role in improving the effectiveness and precision of data annotation. Their ability to automate annotation tasks (Zhang et al., 2022b), ensure consistency across large volumes of data, and adapt through fine-tuning or prompting for specific domains (Song et al., 2023), significantly mitigates the challenges encountered with traditional annotation methods, setting a new standard for what is achievable in the realm of NLP.

### v2

@@ 5.1.2. Evaluation Metrics
For position bias, we use Position Consistency as a metric, which quantifies how frequently a judge model selects the same response after changing their positions. Formally, given $N$ samples $\{(q_{i},r1_{i},r2_{i})\}^{N}_{i=1}$, for each sample $(q_{i},r1_{i},r2_{i})$, we query the LLM evaluator with two prompts $P(q_{i},r1_{i},r2_{i})$ and $P(q_{i},r2_{i},r1_{i})$, and obtain corresponding two evaluation results $S_{i}^{r12}$ and $S_{i}^{r21}$. Each $S_{i}$ is $r1_{i}$, $r2_{i}$ or ”TIE”. Then we calculate the position consistency as follows:
@@ 5.1.3. Target LLMs and Strategies
For LLMs, we selected three LLMs commonly used in automatic evaluation as evaluators, including closed-source LLMs GPT-4, GPT-3.5, and open-source LLMs Qwen2.5-7B, LLaMA3-8B, Mistral-7B, and Mixtral-8$\times$7B.
For improvement strategies, we selected Providing Evaluations with Explanations, Self Validation, Summarize by Multiple Rounds, and Vote by Multiple LLMs, since these strategies are all very straightforward and relatively uniform in many works. We adopt GPT-3.5 as the base evaluator when conducting meta-evaluation for these improvement strategies.
@@ 5.1.4. Model Configuration
For the closed-source LLMs, we use the official API provided by OpenAI to interact with LLMs. The model versions we selected are GPT-4-turbo and GPT-3.5-turbo, specifically referencing gpt-4-turbo-2024-04-09 and gpt-3.5-turbo-0125 respectively [^2] https://platform.openai.com/docs/models.
For the open-source LLM, we adopt Qwen2.5-7B-Instruct [^3] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct, Meta-Llama-3-8B-Instruct [^4] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct, Mistral-7B-Instruct-v0.3555https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3, Mixtral-8$\times$7B-Instruct-v0.1666https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1, deployed on an Ubuntu machine equipped with a 40GB NVIDIA A100 GPU.
To stabilize the evaluation results of LLMs, we set the hyper-parameter temperature to 0 to reduce the impact of randomness in LLMs’ output. For Summarize by Multiple Rounds, we conduct 5 rounds for each sample, and verify the effects of three different processing methods for results of multiple rounds: majority voting(- majority@5), taking the mean score(- mean@5), and taking the best score(- best@5). For Vote by Multiple LLMs, we conduct experiments on two settings, each involving three LLMs. Setting 1 consists of GPT-4-turbo, GPT-3.5-turbo, and LLaMA3-8B-Instruct, while setting 2 consists of GPT-4-turbo, GPT-3.5-turbo, and Qwen2.5-7B-Instruct.
@@ 5.2. Experiment Results and Analysis
LLMs Alignment Biases with Position Length Concre- Empty Content Nested Familiar Human teness Reference Continuation Instruction Knowledge (n=5106) (n=2633) (n=34) (n=28) (n=26) (n=24) (n=24) (n=24) GPT-4-turbo 61.57 80.49 91.18 89.29 65.38 95.83 70.83 100.0 GPT-3.5-turbo 54.72 68.78 20.59 64.29 23.08 91.67 58.33 54.17 Qwen2.5-7B-Instruct 56.54 63.50 64.71 71.43 69.23 91.67 45.83 83.33 LLaMA3-8B-Instruct 50.72 38.85 20.59 57.14 65.38 75.00 45.83 54.17 Mistral-7B-Instruct-v0.3 55.42 59.78 26.47 67.86 53.85 66.67 37.50 41.67 Mixtral-8$\times$7B-Instruct-v0.1 56.29 59.06 50.00 78.57 42.31 83.33 29.17 83.33
@@
Table 2. The meta-evaluation results for different LLMs. All the values are percentages.
Improvement Strategies Alignment Biases with Position Length Concre- Empty Content Nested Familiar Human teness Reference Continuation Instruction Knowledge (n=5106) (n=2633) (n=34) (n=28) (n=26) (n=24) (n=24) (n=24) GPT-3.5-turbo - base 54.72 68.78 20.59 64.29 23.08 91.67 58.33 54.17 - w/ explanation 52.47 48.97 35.29 60.71 38.46 91.67 41.67 50.00 - w/ self-validation 54.86 69.31 23.53 60.71 23.08 91.67 41.67 50.00 - w/ multi rounds     - majority@5 54.68 70 11 26.47 67.86 23.08 95.83 54.17 50.00     - mean@5 54.72 69.58 11.76 57.14 26.92 87.50 50.00 50.00     - best-of-5 51.95 58.72 5.88 42.86 19.23 87.50 37.50 45.83 multi LLMs (set 1) 57.66 32.28 26.47 64.28 46.15 87.50 66.67 62.50 multi LLMs (set 2) 58.19 70.98 64.71 71.43 69.23 91.67 45.83 83.33
Table 3. The meta-evaluation results for different strategies based on GPT-3.5-turbo. All the values are percentages.
5.2.1. Comparison with Different LLMs
The experiment results on different LLMs are shown in Table 2. Comparing the evaluation performance of different LLMs, we found GPT-4 outperformed other LLMs with a large margin across all meta-evaluation dimensions and showed fewer biases. Therefore, when conditions allow, using GPT-4 as an automated evaluator may obtain more objective and less biased evaluation results. For open-source LLMs, we found that Qwen2.5-7B-Instruct showed exceptional evaluation capabilities, outperforming other open-source LLMs in the experiments. Moreover, it surpassed GPT-3.5-turbo in most dimensions except for Position Bias and Nested Instruction Bias, indicating that it can be a promising choice as an open source LLM-as-a-Judge, with the potential to serve as a robust base model for specialized evaluators in specific scenarios.
Additionally, we observed that, apart from Concreteness Bias and Content Continuation Bias, the performance of LLMs except GPT-4-turbo was generally poor, particularly in the Length Bias. And even GPT-4-turbo experienced substantial performance degradation in Empty Reference Bias and Nested Instruction Bias. While Position Bias can be mitigated by swapping the positions of the evaluation contents, addressing other biases may require researchers to explore more effective evaluation strategies. Meanwhile, we also observed that there was not much difference in alignment with human among different LLMs in the experiments, and all of them showed significant room for improvement.
5.2.2. Comparison with Different Strategies
@@
Table 3 shows the effectiveness of different improvement strategies for enhancing the evaluation performance of GPT-3.5-turbo. The results reveals that no all evaluation strategies effectively improve LLM-as-a-judge’s evaluation outcomes. Providing with Explanation (w/ explanation) provides interpretability by offering reasons alongside evaluation scores or selections, which aids in logical backtracking during human review. However, in terms of evaluation performance and bias mitigation, it generally has a negative impact. This performance decline is speculated to be caused by deeper biases introduced by self-explanation. Self Validation (w/ self-validation) shows minimal effectiveness, likely due to the LLMs’ overconfidence, which may limit its re-evaluation efforts during self-validation. We will further discuss this limitation in Section 7.1.
Summarize by Multiple Rounds with majority voting (w/ majority@5) is a strategy with clear benefits, showing improvements across multiple dimensions. It suggests that taking the majority voting results from repeated evaluations help reduce the impact of randomness in LLMs, thereby addressing bias issues. However, Summarize by Multiple Rounds with taking mean score (w/ mean@5) or with taking best score (w/ best-of-5) did not improve the evaluation performance and even had some adverse effects. Compared to w/ majority@5, which selects the major result from multiple rounds, w/ mean@5 might include results with biases in the mean score calculation, and similarly w/ best-of-5 could potentially select overly high scores influenced by biases. Therefore, the latter two strategies do not effectively mitigate the impact of biases on automated evaluation.
The evaluation results of Vote by Multiple LLMs (multi LLMs set 1 and set 2) are closely related to the LLM selection. Comparing set 1 and set 2, where LLaMA3-8B-Instruct was replaced by Qwen2.5-7B-Instruct in set 2, it revealed significant differences in performance across various dimensions. In set 1, the poor performance of GPT-3.5-turbo and LLaMA3-8B-Instruct in the Length Bias negatively impacted the overall performance, whereas in set 2, the performance in this dimension was better, which was aligned with Qwen2.5-7B-Instruct. Similar trends were observed in dimenstions like Position Bias, Familiar Knowledge Bias, and so on. This suggests that when multiple LLMs are adopted for joint evaluation, the differences between their evaluation performances must be carefully considered.
5.2.3. Summary
@@
Due to the inherent capabilities and potential risks of LLMs, common improvement strategies for LLM-as-a-judge are not fully effective in improving the evaluation performance or mitigating biases. The limitations and challenges will be further discussed in Section 7.
Based on the current experimental analysis, an empirical strategy for pairwise comparison evaluation tasks is to select more powerful LLMs and to adopt two evaluation strategies: one is swapping the positions of the evaluation contents, the other is taking the majority voting results from multiple rounds of evaluation, which can effectively mitigate biases. As for improving the alignment with humans, further exploration is still needed.
@@ 6.2.1. Finance
LLMs have demonstrated significant potential in the finance domain, particularly in tasks such as forecasting, anomaly detection, and personalized text generation (Zhao et al., 2024a), thereby driving an increasing demand for LLM evaluators.
In the context of LLM-as-a-judge applications within finance, expert knowledge is crucial for domain-specific evaluations. Current research can be divided into two areas: one focuses on designing LLM-based evaluators that leverage expert knowledge for specific tasks. For instance, Brief et al. (2024) conducted a case study on multi-task fine-tuning in finance to enhance LLM performance (Brief et al., 2024), while Yu et al. (2024) introduced FinCon, a multi-agent system that uses conceptual verbal reinforcement to improve financial decision-making (Yu et al., 2024a). The second area of research aims to provide benchmarks to evaluate and enhance LLMs’ understanding of domain-specific knowledge. These benchmarks include user-feedback-based UCFE (Yang et al., 2024), IndoCareer—a dataset of professional exam questions (Koto, 2024), and AI-generated domain-specific evaluation sets (Raju et al., 2024).
Additionally, the concept of LLM-as-a-judge shows promising applications in credit scoring (Yoon, 2023; Babaei and Giudici, 2024) and Environmental, Social, and Governance (ESG) scoring (Zhao et al., 2024a). This work remains in its early stages, necessitating further exploration to refine evaluation methods and expand applications in the finance domain.
@@ 6.2.2. Law
The second category of research involves creating benchmarks for evaluating LLM applicability in legal scenarios. Examples include multi-domain evaluation sets, such as the IndoCareer dataset for professional exams in Indonesia (Koto, 2024) and LegalBench, a collaboratively built benchmark for assessing legal reasoning capabilities in LLMs across multiple domains and languages (Guha et al., 2023). These benchmarks are often language-specific like LexEval for Chinese legal texts (Li et al., 2024a) and Eval-RAG for Korean  (Ryu et al., 2023). Other benchmarks target specific attributes, such as ethics (Zhang et al., 2024a) and harmfulness (Andriushchenko et al., 2024).
@@ 6.2.4. Others
LLMs have also been employed as evaluators to enhance efficiency and consistency across various fields. In software engineering, a method was proposed for using LLMs to evaluate bug report summarizations, demonstrating high accuracy in assessing correctness and completeness, even surpassing human evaluators who experienced fatigue (Kumar et al., 2024a). This approach offers a scalable solution for evaluation. In education, automated essay scoring and revising were explored using open-source LLMs, achieving performance comparable to traditional deep-learning models. Techniques such as few-shot learning and prompt tuning improved scoring accuracy, while revisions effectively enhanced essay quality without compromising original meaning (Song et al., 2024b). In content moderation, an LLM-based approach was developed to identify rule violations on platforms like Reddit, achieving high true-negative rates but encountering challenges with complex rule interpretation, emphasizing the necessity of human oversight for nuanced cases (Kolla et al., 2024). In behavioral sciences, the LLM-as-a-Judge framework was evaluated for assessing user preferences based on personas, revealing limitations in reliability and consistency due to oversimplified personas, but improved significantly through verbal uncertainty estimation, achieving high agreement with human evaluations for high-certainty cases (Dong et al., 2024). These applications of LLMs as evaluators highlight their growing potential in diverse sectors, emphasizing the need for integrating domain-specific knowledge and refining methodologies.
@@ 7.1. Reliability
Evaluating the reliability of LLMs when used as judges reveals several pressing challenges. Both human and LLM judges exhibit biases, which raises concerns regarding the consistency and fairness of their evaluations. Specifically, human judges are also found to have inherent bias (Zheng et al., 2023b; Wu and Aji, 2023) and may not even provide reliable answers (Clark et al., 2021; Hämäläinen et al., 2023). As an alternative to humans, LLM evaluations are also found to have certain biases, and the annotation results require more evaluation (Pangakis et al., 2023), as we discussed in § 4. The bias of LLM-as-a-judge is more due to the fact that LLM is a probabilistic model, as we have defined in § 4. Moreover, Reinforcement Learning with Human Feedback (RLHF) improves LLM performance by aligning them with human preferences. However, ensuring models trained with RLHF (Lambert et al., 2024) produce robust and consistent outputs remains an ongoing challenge.
@@
Instruction-tuned LLMs have been demonstrated to possess the issue of overconfidence, which means they tend to offer overly favorable scores when evaluating their own responses (Tian et al., 2023). The overconfidence is also highly likely to exist in the scenario of LLM-as-a-judge, which is also engaged in evaluating the responses generated by LLMs. Consequently, when LLM-as-a-judge is utilized with the latest LLMs, which are typically instruction-tuned, the existence and impact of overconfidence need to be meticulously examined.
@@ 8.1. More Reliable LLM-as-a-Judge
inally, the uncertain and evolving nature of robustness risks underscores the necessity of proactive mitigation strategies. These strategies should include the development of adversarial training techniques tailored to judgment tasks, the integration of robust uncertainty quantification methods, and the implementation of human-in-the-loop systems to oversee critical decisions. By addressing these challenges, we can build more resilient and dependable systems capable of maintaining high levels of reliability even under adversarial conditions.
@@ 8.2. LLM-as-a-Judge for Data Annotation
Despite its wide applications, data annotation poses significant challenges for current machine learning models due to the complexity, subjectivity, and diversity of data. This process requires domain expertise and is resource-intensive, particularly when manually labeling large datasets. Advanced LLMs such as GPT-4 (OpenAI, 2023a), Gemini (Google, 2023), and LLaMA-2 (Touvron et al., 2023c) offer a promising opportunity to revolutionize data annotation. LLMs serve as more than just tools but play a crucial role in improving the effectiveness and precision of data annotation. Their ability to automate annotation tasks (Zhang et al., 2023e), ensure consistency across large volumes of data, and adapt through fine-tuning or prompting for specific domains (Song et al., 2024a), significantly mitigates the challenges encountered with traditional annotation methods, setting a new standard for what is achievable in the realm of NLP.